OpenStudy (anonymous):

the product of two consecutive positive even number is 1520. what are the numbers? (enter solutions from smallest to largest)

OpenStudy (lgbasallote):

let x = one number x + 2= next even number the product is 1520 so x(x+2) = 1520 x^2 + 2x = 1520 x^2 - 2x - 1520 = 0 you're free to use quadratic equation to solve for x now :)